OpenStudy (anonymous):

I need help factoring: 2x^2+13x+15

OpenStudy (anonymous):

x(2x+13) + 15

OpenStudy (anonymous):

(x+5)(2x+3)

OpenStudy (anonymous):

so then, what would be the zeros? The orginal question was to find all the zeros of: 2x^3-3x^2-50x+75

jimthompson5910 (jim_thompson5910):

Multiply the first coefficient 2 and the last term 15 to get 2*15 = 30 Then you need to find two numbers that multiply to 30 and add to 13 (the middle coefficient) These two numbers are 10 and 3 since 10*3 = 30 and 10+3 = 13

jimthompson5910 (jim_thompson5910):

so that means we break up 13x into 10x+3x and factor by grouping

OpenStudy (anonymous):

-5, -3/2

jimthompson5910 (jim_thompson5910):

2x^2+13x+15 2x^2+10x+3x+15 (2x^2+10x)+(3x+15) 2x(x+5)+(3x+15) 2x(x+5)+3(x+5) (2x+3)(x+5)

OpenStudy (anonymous):

I figured it out, Thanks!!!

jimthompson5910 (jim_thompson5910):

If you want to factor 2x^3-3x^2-50x+75, then... 2x^3-3x^2-50x+75 (2x^3-3x^2)+(-50x+75) x^2(2x-3)+(-50x+75) x^2(2x-3)-25(2x-3) (x^2-25)(2x-3) (x-5)(x+5)(2x-3) ... use the difference of squares factoring rule So that shows us that 2x^3-3x^2-50x+75 factors to (x-5)(x+5)(2x-3)

jimthompson5910 (jim_thompson5910):

Ok I'm glad you figured it all out